OpenStudy (anonymous):

In a circle, a central angle that measures 36° intercepts arc AC. What is the measure of arc AC?

OpenStudy (anonymous):

radius is not given so i think solution of this question is not possible

OpenStudy (anonymous):

MAybe they want it in terms of r: \[Circumference = 2 \pi r\] \[2 \pi r * 36/360 = \pi r/5\]

OpenStudy (anonymous):

express 37 in radian that is 37xpi/180=ac/r hence AC=37xpixr?180